Sandra is Y years old. In 14 years her age will be three time her age fourteen years ago.
Answer by Maths68(1474)

Sandra = y years old
In 14 years
Sandra will be = y+14 years old
14 years ago
Sandra was = y-14
In 14 years her age will be three time her age fourteen years ago
3(y-14)=y+14
3y-42=y+14
3y-y=14+42
2y=56
2y/2=56/2
y=28

Sandra = 28 years old